# ScrollTimes: Tracing the Provenance of Paintings as a Window into History

### v1

@@ Abstract
Digital humanities research has flourished due to the diverse artifacts available in cultural heritage databases. However, over-reliance on a single artifact type can result in poor contextualization and a constrained understanding of historical context. We collaborated with art historians to examine handscrolls, a form of traditional Chinese painting which offers a wealth of data for historical analysis and provides a unique opportunity for understanding history through artwork. We propose ScrollTimes, a visual analysis system for tracing handscroll historic context by linking multiple data sources. Specifically, a unique layout is developed for efficiently viewing long handscrolls. Using image processing techniques and language models, we extract, verify, and supplement elements in handscrolls with different cultural heritage databases. Furthermore, interactive biographies are constructed for handscrolls to uncover their historical narratives, provenance trajectories, and artistic legacies. Validated through case studies and expert interviews, our approach offers a window into history, fostering a holistic understanding of handscroll provenance and historical significance.
@@
traditional Chinese painting, handscroll, painting biography, neural networks, seam carving
@@
The visual analysis system ScrollTimes enables art historians to analyze handscrolls by combining multiple databases and constructing biographies of paintings that fit their research interests. In the case of exploring the renowned handscroll “Autumn Colors on the Que and Hua Mountains,” historians first utilized the Handscroll Features View (A) to understand the handscroll elements, such as the core painting, seals, and inscriptions. Then, they further explored the elements of interest in the Seals and Inscriptions View (B) and validated the information using the Linked Data View (B4). Finally, they constructed a biography of the painting that matched their research interests using the Biography View (C).
@@ 1 Introduction
The emergence of cultural heritage databases has significantly propelled the development of Digital Humanities (DH). This advancement has transformed traditional research workflows by streamlining manual data collection and annotation, thus enabling scholars to analyze vast amounts of data more effectively [65, 18]. Various DH studies have employed cultural heritage databases that collect only a single artifact type (e.g., ancient literature [69, 68], paintings [39], and sculptures [66]). Although different types of cultural artifacts inherently originate from specific historical contexts and often recount the same historical events, they exhibit diverse themes, formats, meanings, and purposes. Relying on only one artifact type results in a restricted understanding of history [27]. For example, computational painting analysis tends to focus more on the aesthetic value of the painting itself (e.g., research on the style [39], patterns [26], and creation processes [59]), while largely neglecting the painting’s historical value as an artifact.
In this study, we explore the “handscroll” [45], a distinguished format in traditional Chinese painting, as the primary artifact type to reveal the historical context embedded in these artworks, together with multiple data sources. The handscroll is selected because of its unique creation process, which serves as a rich source of historical information over a long period of time. The core painting (Fig. 1B, usually landscapes or narrative stories) is typically drawn on a long piece of silk or paper from right to left. This wide aspect ratio enables the painter to depict the flow of time and the variations in space upon a single handscroll. Upon the completion of the core painting, the painter often adds inscriptions (Fig. 1A, text describing the motivation behind the creation) and seals (Fig. 1C, vermillion imprint indicating their signature). At this point, a handscroll is “born”. However, its development does not cease, as the “life” of the handscroll continues. As future generations appreciate the handscroll, they frequently contribute additional content, such as mounting a new frontispiece or colophon with their own inscriptions (including critiques or historical event records) and seals (indicating appreciation). Therefore, the handscroll not only records history but also becomes a part of it [28].
@@
The centuries-old handscrolls resonate with historic significance, and art historians have been studying their provenance over time based on their rich components. We propose creating biographies for handscrolls to explore their historical narratives. By recording their “life stories” (i.e., appreciation records and anecdotes) and exploring the events involving figures who interacted with the paintings, we aim to use the paintings to verify historical events, providing new perspectives and insights into the longitudinal study. In addition, since traditional Chinese painting places great emphasis on copying and transferring styles [16], the inheritance of painting styles can be traced through comparisons of similarities, akin to viewing the “genealogy” of a painting (i.e., other paintings that were influenced by it). The relationships between painters can also be discovered by linking with other cultural heritage databases, thereby offering cross-sectional data to broaden the scope of historical research.
In collaboration with art historians, we have identified three primary challenges in building biographies of handscrolls: 1) the difficulty in viewing handscrolls due to their wide aspect ratios, 2) the inefficiency of conventional methods in extracting and analyzing inscriptions and seals to understand the provenance of paintings, and 3) the overwhelming task of aggregating multiple data sources, which relies on serendipity [65] if prior knowledge is not incorporated into the workflow.
To enhance the viewing experience, we leverage seam carving [9] to reduce the white space in handscrolls. Then the wide and narrow paintings are transformed into a circular format, enabling more efficient, interactive viewing for art historians. To capture essential components, we employ deep neural networks [14, 31, 32] for extracting core painting and seals, and then retrieving implicit information from multiple databases. Moreover, we utilize an NLP model [63] to extract entities such as time, location, figures, and things from inscriptions in Classical Chinese, then multiple databases are used to verify existing data and supplement missing data. Finally, we propose the ScrollTimes, an interactive visual analysis system to integrate multiple databases and offer validation and supplementation from various perspectives, assisting art historians in comprehending painting biographies and analyzing the provenance history of the paintings. In summary, this study has the following contributions:
  • We developed a new painting viewing method, the circular view, based on handscroll characteristics, enabling art historians to swiftly view paintings.
  • We used automatic methods and multiple databases to extract, validate, and supplement elements in paintings, allowing art historians to create painting biographies interactively using ScrollTimes.
  • We conducted case studies and expert interviews to demonstrate the effectiveness and usability of our approach.
@@ 2 Related Works
2.1 Visualize Linked Data for Cultural Heritage
With the advent of digital cultural heritage databases around the world, open linked data provides a formidable foundation for reconstructing the historical evolution of culture [64, 65, 29, 52, 21]. Researchers leverage various visual analysis techniques such as Semantic Web technology [19, 43, 24, 29], multi-temporal and spatial reconstruction [53, 12], and web-GIS [50] to structure and elegantly display the heterogeneous data. These works largely promote resource sharing and comprehensive information exploration.
Researchers are increasingly cognizant of the significance of cultural heritage databases [24, 25, 65, 54]. For instance, some studies use semantic web technology and natural language processing methods to enhance access to cultural information and address the public challenges of managing small collections. Meanwhile, visualization methods for digital cultural heritage collections have been scrutinized and reviewed, and new applications (such as “Casual Learn” [57]) and tools (such as CHEST [30] and “InTaVia” [48]) have been introduced. The “InTaVia” project has developed a comprehensive visualization approach for a multi-modal historical knowledge graph while emphasizing the importance of a user-centered design process. Nevertheless, the credibility of the visual analysis results for cultural heritage remains a challenge due to the intricacy and diversity of cultural heritage data, which often engenders ambiguity and misunderstanding. Additionally, current platforms lack better integration of prior domain expert knowledge, which plays a crucial role in professional usages.
Our system endeavors to furnish art historians with validated entity information from multiple data sources, integrates the information for art historians to make inferences with prior knowledge, and ultimately offers easy-to-use interactions to refine the information based on requirements.
@@ 2.2 Computational Approaches for Analyzing Painting
Recently, a surge of interest has emerged in constructing painting databases from various angles, including the painter [42], theme [46, 37], and style [40] of paintings. Some of them support cross-modal retrieval by combining text and image features [23, 47]. These databases have greatly supported the interpretation [27, 26], creation [61, 67], and exhibition [36, 41] of traditional Chinese painting. However, these works tend to focus more on the aesthetic value of the painting itself. MTFFNet [39] has been proposed to analyze the painting style of painters, Fanet al. [26] focus on discovering painting patterns, and Tanget al. [59] reproduce the painting process for better painting education. The historical value of paintings remains unexplored.
@@
This study aims to explore the historical value of Chinese handscroll paintings and develop a visual analysis system tailored to the needs of art historians. To this end, we employ automated methods to extract handscroll elements and validate them by cross-referencing with relevant historical contexts from external cultural heritage databases. We then construct biographies for the paintings and utilize event sequence visualization techniques [33, 44] to comprehensively demonstrate the historical significance.
@@ 3 Background
This section provides an overview of the domain background and outlines the requirements obtained through in-depth interviews with domain experts. Over the past year, we have collaborated closely with five highly experienced art historians. This group includes three professors (A1-A3), each possessing over twenty years of research experience in the field of traditional Chinese painting history, and two Ph.D. students (A4 and A5) with nine years of practical experience in traditional Chinese painting. In addition, we seek guidance on cultural heritage data processing from two editors (E1 and E2) of the China Biographical DataBase [11]. Our collaboration took place in three stages: task analysis (Section 3.2), system design (Section 5), and system evaluation (Section 6). We first present the data description in Section 3.1.
@@ 3.1 Data Description
This study collects 1380 handscroll data (HD) from the open dataset platform [49]. Each handscroll contains the following data elements:
@@
However, these data elements are usually ambiguous and incomplete, which hinders art historians’ systematic analysis. To supplement the handscroll data and perform cross-validation, we selectively collect data from five databases containing information on Chinese historical figures, time periods, locations, and events. These databases (Fig. 2) are chosen based on their relevance to the study of Chinese painting history and their reputation within the academic community. The selected databases include CBDB, CHFSD (Chinese Historical Figure Seal Data) [2], PerAD (Person Authority Database) [5], PlaAD (Place Authority Database) [6], and TAD (Time Authority Database) [7].
@@
3.2 Requirement and Task Analysis
We interviewed our collaborators to gain insights into their research process and inform our system design. Accordingly, we identified the system requirements and design tasks necessary to address the challenges that art historians have been facing when performing conventional analysis using current techniques.
R1: Efficiently browse handscroll and capture its features. Handscrolls’ wide aspect ratio makes traditional study methods challenging, as art historians need to repeatedly scan the handscroll to connect the core painting, inscriptions, and seals to gain insights. A5 indicates that this may take up to 10 hours. Our system should provide an efficient solution for quickly browsing and capturing handscroll features.
  1. T1
    Provide interactive browsing methods to optimize the viewing of handscrolls. Digitization of cultural artifacts offers coarse-to-fine viewing of painting details. However, handscrolls’ wide aspect ratios and excessive white spaces pose challenges to traditional viewing methods. The handscroll viewing experience shall be optimized with convenient and interactive browsing methods.
  2. T2
    Extract and visualize the features of handscrolls. Handscrolls contain elements with rich information, and there exist latent correlations among these elements. However, manual acquisition and connection are time-consuming. To improve the efficiency of handscroll analysis, we intend to utilize computer technology to extract features and visualize the correlations.
R2: Validate and supplement handscroll features from multiple perspectives automatically. To enhance the credibility of analysis results, art historians habitually cross-reference a handscroll’s elements from multiple perspectives to reduce ambiguity and incompleteness. And additional historical context is required to attain a comprehensive understanding of the handscroll. Our system should offer an efficient and effective approach to releasing historians from heavy labor.
  1. T3
    Cross-reference handscroll features from multiple databases. A1 points out that the traditional analysis workflow often necessitates perusing extensive literature across multiple platforms to cross-reference handscroll elements, which is time-consuming. It is imperative to establish automated methods that employ various databases for cross-references handscroll elements, improving the reliability of art historians’ analysis.
  2. T4
    Provide rich historical contextual information for handscroll features. Art historians not only need to connect the elements within the artwork but also relate to external information, which can provide them with new research perspectives and insights. However, relying on a single data source cannot achieve this. We need to combine different cultural heritage databases to provide rich historical contextual information.
R3: Construct painting biographies according to user’s requirements interactively. Art historians have varying standards for constructing biographies of paintings based on their own research interests. For example, A4 is skilled at creating rigorous and completed painting biographies primarily from a literary perspective, while A5 hopes to freely explore the biographies based on the similarity of core paintings. After automatically generating the painting biographies, our system should further enable art historians to customize the painting biographies from different perspectives in an interactive manner.
  1. T5
    Refine painting biographies interactively. Historical data in databases is often limited by data sources, e.g., rare literary documents, which can result in missing or incomplete information in painting biographies. Interactive supplementation by art historians can effectively address this issue. The system provides data from multiple sources, and art historians can use their prior knowledge to integrate and supplement the data elements. This interaction can help to create painting biographies that are comprehensive and more aligned with their interests.
  2. T6
    Adjust painting biographies accordingly. A1 and A3 indicate that art historians have different interests in the information level of painting biographies. Typically, the content within the painting itself, the complementary information from external data sources, and the cross-data recommendations from automatic computer technologies are the three levels that art historians are mostly concerned with. Therefore, art historians should be enabled to interactively adjust different information levels of the painting biography accordingly.
4 Model
This section introduces the techniques utilized for information extraction and processing of various elements in the handscroll, including inscriptions, core painting, and seals. Furthermore, it elaborates on the rigorous process of verifying and supplementing the extracted information with multiple databases to obtain a comprehensive understanding of the events depicted in the handscroll.
4.1 Handscroll Feature Extraction
We commence our approach by leveraging image processing and natural language processing techniques to extract pertinent information from the handscroll (T1).

### v2

@@ Abstract
The study of cultural artifact provenance, tracing ownership and preservation, holds significant importance in archaeology and art history. Modern technology has advanced this field, yet challenges persist, including recognizing evidence from diverse sources, integrating sociocultural context, and enhancing interactive automation for comprehensive provenance analysis. In collaboration with art historians, we examined the handscroll, a traditional Chinese painting form that provides a rich source of historical data and a unique opportunity to explore history through cultural artifacts. We present a three-tiered methodology encompassing artifact, contextual, and provenance levels, designed to create a “Biography” for handscroll. Our approach incorporates the application of image processing techniques and language models to extract, validate, and augment elements within handscroll using various cultural heritage databases. To facilitate efficient analysis of non-contiguous extracted elements, we have developed a distinctive layout. Additionally, we introduce ScrollTimes, a visual analysis system tailored to support the three-tiered analysis of handscroll, allowing art historians to interactively create biographies tailored to their interests. Validated through case studies and expert interviews, our approach offers a window into history, fostering a holistic understanding of handscroll provenance and historical significance.
@@
Visual analytics, Digital humanities, Painting analysis, Traditional Chinese painting
@@
ScrollTimes enables art historians to analyze handscrolls by combining multiple databases and constructing biographies of paintings that fit their research interests. In the case of exploring the renowned handscroll “Autumn Colors on the Que and Hua Mountains,” historians first utilized (A) the Handscroll Features View to understand the handscroll elements, such as the core painting, seals, and inscriptions. Then, they further explored the elements of interest in (A3) the Seals and Inscriptions View and validated the information using (B) the Linked Data View. Finally, they constructed a biography of the painting that matched their research interests using (C) the Painting Biography View, (D) the Similar Painting View and (E) the Uncertainty Data View.
@@ 1 Introduction
In archaeology and art history, the provenance of cultural artifacts–tracing their ownership and preservation–has been a prominent research focus [26]. Such provenance not only manifests an artifact’s authenticity, but also boost its valuation by associating the legacies of previous owners [60]. Moreover, uncovering provenance for cultural artifacts can transcend them from mere static collectibles to vibrant pieces of history, offering new insights and narratives. With the convenient access to cultural artifact databases [20, 58, 59], digitalized copies of cultural artifacts and their related information have been used to analyze their provenance. For instance, “Mapping painting” [4] and “Hieronymus Bosch” [3] visualized the provenance of paintings to engage the wider public in artifacts’ preservation efforts.
However, these projects are insufficient in addressing specific challenges faced by professionals in in-depth provenance analysis [49]. First, the provenance of cultural artifacts can be documented in other sources and directly on the artifacts. For instance, texture analysis reveals the evolution of style, and inscriptions on the artifacts record changes in ownership. However, recognizing these evidences currently rely on domain expertise, hindering the efficiency of provenance analysis. Second, each artifact captures the historical context of its era, influenced by its successive owners’ narratives. Current research often overlooks the rich sociocultural background against which these cultural artifacts carry during their transmission. Consequently, this calls for an integrative approach to connect multiple databases to uncover the intertwined stories of creators, owners, and the historical events they were involved. Lastly, there is a pressing need to develop sophisticated interactive systems that emphasize automation. These systems should enable researchers to efficiently construct in-depth provenance narratives. Currently, the visual storytelling tools for “biographies of things” [49] are limited and often demand excessive manual input.
@@
In this study, we explore the provenance analysis on “handscroll” [51], a distinguished format in traditional Chinese painting. We aim to reveal the historical context embedded in these artworks by combining multiple data sources. The handscroll is selected because of its unique creation process, which provides rich historical information over a long period of time. The core painting (Fig. 1B, usually landscapes or narrative stories) is typically drawn on a long piece of silk or paper from right to left. This wide aspect ratio enables the painter to depict the flow of time and the variations in space upon a single handscroll, which can be considered as a scrollytelling. Upon the completion of the core painting, the painter often adds inscriptions (Fig. 1A, text describing the motivation behind the creation) and seals (Fig. 1C, vermillion imprint indicating their signature). At this point, a handscroll is “born.” However, its development does not cease, as the “life” of the handscroll continues. When future generations acquire and appreciate the handscroll, they frequently contribute additional content, such as mounting a new frontispiece or colophon with their own inscriptions (including critiques or historical event records) and seals (indicating appreciation). Therefore, the handscroll not only records history but also becomes a part of it [30].
In collaboration with art historians, we propose creating “Biography” for handscroll to explore their historical narratives. It encompasses three distinct levels:
  • At the artifact level, it describes the characteristics and appearance of cultural artifacts. To improve the efficiency of summarizing the characteristics of the handscroll, we automatically extract the elements embedded within the handscroll. Considering the unique appearance of handscroll, we implement seam-carving [9] method to optimize layouts facilitating art historians in the focused analysis of non-adjacent extracted elements.
  • At the contextual level, it constructs a social network related to cultural artifacts. We integrate multiple databases to verify extracted elements and supplement information, assisting art historians in uncovering the relationships among historical figures and events related to the handscroll.
  • At the provenance level, it chronicles the life experiences of cultural artifacts. We combine Artifact and Contextual information through temporal visualization. Additionally, we employ a hierarchical algorithm to rank the significance of historical figures, allowing art historians to intuitively observe the ups and downs of handscroll’s life experiences and compare the features of handscroll provenance across different time periods.
Based on three levels, we propose ScrollTimes, an interactive visual analytic system that explores the provenance of cultural artefacts. It automatically generates handscroll biographies and assists art historians in customizing them. We evaluated the system’s effectiveness and usability through two case studies and expert interviews. Finally, we have created 1380 handscroll biographies and open-sourced the system at https://scrolltimes.cn.
@@ 2 Related Works
2.1 Visualizing Linked Data for Cultural Heritage
With the advent of digital cultural heritage databases, open linked data provides a solid foundation for reconstructing the historical evolution of culture [76, 19, 53]. Researchers leverage various computational techniques such as semantic web technology [17, 23], multi-temporal and spatial reconstruction [63, 11], and web-GIS [61] to structure and elegantly display the heterogeneous data. These works largely promote resource sharing and comprehensive information exploration.
Researchers have increasingly integrated multiple cultural heritage databases to discover historical insight [81], enhance access to cultural information [64], and manage multiple small collections [24]. Meanwhile, visualization methods for digital cultural heritage collections have been gaining popularity [76], and new applications (e.g., Casual Learn [67]) and tools (e.g., CHEST [31] and InTaVia [57]) have been introduced. Visual analytics approach has the advantage of contextualizing multi-modal historical knowledge while emphasizing the importance of a user-centered design process. A variety of heritage items has been explored in depth, including ancient books [35], paintings [71], and cultural artifacts [78, 52]. In this paper, we focus on paintings in the form of handscroll, which has high cultural values and multiple references in different sources (e.g., poems and documentation).
Nevertheless, the credibility of the visual analysis results for cultural heritage remains a challenge due to the intricacy and diversity of cultural heritage data, which often engenders ambiguity and misunderstanding [83]. Additionally, current platforms lack better integration of prior domain expert knowledge, which plays a crucial role in professional usages. We improve over existing works in terms of providing art historians with validated entity information from multiple data sources and integrating this information to enable them to make informed inferences based on their prior knowledge.
@@ 2.2 Computational Approaches for Analyzing Painting
Recently, a surge of interest has emerged in constructing painting databases from various angles, including the painter [47], theme [55, 41, 80], styles [44], and even AI-generated paintings [28]. Some support cross-modal retrieval by combining text and image features [22, 56]. These databases have greatly supported the interpretation [27, 25], creation [74, 79], and exhibition [40, 45] of traditional Chinese painting [83]. However, they tend to focus more on the aesthetic value of the painting itself. MTFFNet [43] has been proposed to analyze the painters’ painting style. Fanet al. [25] focus on discovering painting patterns. Tanget al. [70] reproduce the painting process for better painting education. The historical value of paintings remains underexplored.
@@
This study aims to explore the historical significance of Chinese handscroll paintings. Different from existing works, we utilize automated methods to extract elements from the paintings and validate them by cross-referencing external cultural heritage databases. The relevant historical contexts can help reveal their historical values. We construct biographies for these paintings and employ event sequence visualization techniques [34, 50, 68], especially in multi-modality [54, 46, 38], to provide a comprehensive portrayal of their historical significance.
@@ 3 Background
This section introduces the domain background and outlines the requirements obtained through in-depth interviews with domain experts.
3.1 Cooperating with Art Historians
Over the past year, we have engaged in a close collaboration with six highly experienced art historians. This group comprises four professors (A1-A4), each boasting more than twenty years of research experience, and two Ph.D. students (A5 and A6) with nine years of practical expertise. A1 and A2 specialize in the study of painting during the Song and Ming Dynasties. A3 and A6 concentrate on stylistic analysis spanning entire generations, while A4 focuses on the analysis of Zen-themed paintings, and A5 specializes in the analysis of paintings with an “Elegant Gathering” theme. The collaborative process can be divided into three key stages as follows:
  • Requirement Summary and Task Identification. We obtained insights into the traditional process of cultural artifact provenance analysis through expert interviews, refining requirements and clarifying system design tasks (Section 3.2).
  • System Design and Enhancement. Experts actively participated in system design, and we iteratively adjusted and improved the system based on their suggestions (Section 4, Section 5).
  • Case Studies and Feedback. Experts conducted case studies aligned with their research interests, providing valuable feedback (Section 6).
3.2 Requirements and Task Analysis
Through interviews with art historians, we have extracted the traditional process of create biography for artifacts, which typically involves formal analysis, contextual analysis, and critical thinking. These experts have emphasized that this process is both labor-intensive and time-consuming, especially in the first two phases. We have streamlined the requirements and design tasks into three levels as bellow.
R1: At the artifact level, analyze handscroll characteristics efficiently. Experts emphasize that the first step in analyzing handscrolls is to conduct a formal analysis of the rich visual elements they contain. Extracting and summarizing these elements aids in uncovering hidden transmission information within the artifact.
  1. T1
    Extracting handscroll elements efficiently. Analyzing handscrolls’ visual properties relies on element extraction. Traditional methods, as indicated by A1-A6, are time-consuming, especially for handscrolls like “Autumn Colors on the Que and Hua Mountains,” with 168 seals and 36 inscriptions. The system needs an automatic method for identifying and extracting elements.
  2. T2
    Visualizing element correlations. A3 and A4 indicate that seal impressions and inscriptions are often scattered along the handscroll. However, the wide aspect ratio of handscrolls hampers efficient analysis, especially when dealing with non-adjacent elements. The system should optimize the handscroll layout to help experts establish element connections quickly.
R2: At the contextual level, establish handscroll social network. A1 and A3 emphasize that artifacts reflect and influence societal and cultural dynamics. Exploring the interactions among artifacts, historical figures, and events deepens our understanding of their historical and cultural significance.
  1. T3
    Identifying historical figures associated with handscroll. Various historical figures are linked to handscrolls, including painters, sealers, inscribers, historical figures mentioned in inscriptions, and those with similar artistic styles. Automated methods are required to identify these historical figures. Additionally, A5 highlights the potential for historical figures to have nickname, requiring the integration of multiple databases for data validation.
  2. T4
    Supplementing contextual information. Historical events related to the identified historical figures are central in constructing social network, aiding experts in understanding their relationship in historical context. The system needs to integrate related databases for comprehensive historical context.
R3: At the provenance level, interactively construct biographies. Experts have indicated the need for a convenient platform to construct handscroll biographies and analyze their transmission characteristics. Simultaneously, they also require multiple analytical dimensions to cater to diverse research demands.
  1. T5
    Constructing handscroll biographies automatically. Since handscrolls often survive for centuries, art historians need to study their cultural significance at different historical moments. The system needs to provide an automated method to link data between the artifact and contextual level in a chronological manner to construct biographies.
  2. T6
    Facilitating interactive exploration based on research interests. Art historians use varying criteria for creating biographies. For example, A3 focuses on stylistic transmission, while A5 emphasizes thematic similarity. They are collectively concerned with the presence of uncertain information within paintings, such as the appearance of historical figures in artwork that are absent from historical records. The system should enable interactive customization with different perspectives.
4 Models
In this section, we initially describe the databases we utilized. Then we have developed automated methods for handscroll elements processing and context information processing (Section 3.2).
@@ 4.1 Data Description
This study collects 1380 handscroll data from the open dataset platform [59]. Each handscroll contains the following data elements:
@@
However, these data elements are usually ambiguous and incomplete, which hinders art historians’ deep analysis. To build the handscroll database and perform cross-validation, we selectively collect data from five databases containing information on Chinese historical figures, time periods, locations, and events. These databases (Fig. 2) are chosen based on their relevance to the study of Chinese painting history and their reputation within the academic community. The selected databases include CBDB, CHFSD (Chinese Historical Figure Seal Data) [2], PerAD (Person Authority Database) [5], PlaAD (Place Authority Database) [6], and TAD (Time Authority Database) [7].
@@
4.2 Handscroll Elements Extraction
At the artifact level, extracting handscroll elements enhances visual properties (R1). We use a natural language model for inscriptions and image processing for core paintings and seals (T1, T3).